OpenStudy (anonymous):

use newton's method with the specified initial approximation x1 to find x3, the thired approximation to the root of thr given equation. (give your answer to four places) x^7+4=0, x1=-1

OpenStudy (dumbcow):

\[x_{n+1} = x_{n} -\frac{f(x)}{f'(x)}\] \[f'(x)=7x^{6}\] \[x_2 = (-1) - \frac{(-1)^{7}+4}{7(-1)^{6}}\] ...

OpenStudy (anonymous):

\[y'=7x^6\] for n=1 \[x2=\frac{ -1-((-1^7)+4) }{ 7(-1^6) }=\frac{ 4 }{ 7 }=0.571\] \[x_{3}=\frac{ 4 }{ 7 }-\frac{ ((\frac{ 4 }{ 7 }) +4)}{ 7(\frac{ 4^6 }{ 7 } )}=0.03464\]

OpenStudy (anonymous):

thats what i did but not sure if its correct or not

OpenStudy (dumbcow):

no you miscalculated....x_2 = -10/7

OpenStudy (anonymous):

how did you got (-10)?

OpenStudy (phi):

the formula, as dumb cow posted is \[ x_{n+1} = x_{n} -\frac{f(x)}{f'(x)} \] notice that you should replace xn with -1 the first time: \[ x_{n+1} = -1 -\frac{f(x)}{f'(x)} \] putting in f(x)=x^7+4 and f'(x)= 7x^6 \[ x_{n+1} = -1 -\frac{x^7+4}{7x^6} \] we must replace x with -1. we get \[ x_{n+1} = -1 -\frac{-1+4}{7} \\ x_{n+1} = -1 -\frac{3}{7} \\ x_{n+1} = -\frac{7}{7} -\frac{3}{7}\\ x_{n+1} = \frac{-7-3}{7} = -\frac{10}{7} \]

OpenStudy (anonymous):

oh okay got it. so the next step will be x= -10/7

OpenStudy (anonymous):

\[x_3=\frac{ -10 }{ 7 }-\frac{ (\frac{ -10 }{ 7 })^7+4}{ 7+(\frac{ -10 }{ 7 })^6 }\] \[=(\frac{ -10 }{ 7 })-(\frac{ -8.142656789 }{ 59.49901827 })\] =-1.2917178 is that correct? Plus is this the final answer?

OpenStudy (dumbcow):

you are correct and yes since they want x_3, that is final answer

OpenStudy (anonymous):

thanks!